# Multi-Scale Fusion for Object Representation

### v2

@@ 4.2 VAE Guidance
To investigate how MSF fuses multi-scale information into VAE guidance, we analyze object separability (Stanic et al., 2024; Lowe et al., 2024). Empirically, better object separability in VAE guidance contributes to better OCL performance. We cluster on the VAE guidance $\bm{X}_{n=1}$ then calculate inter-/intra-cluster distances; We also visualize (Caron et al., 2021) object separability of scale-invariant representations $\{\dot{\bm{X}}_{n}^{\mathrm{si}}\}$ and scale-variant representations $\{\bm{X}_{n}^{\mathrm{sv}}\}$, as well as representations before inter-scale fusion $\{\bm{X}_{n}^{\mathrm{si}}\}$ and representations after intra-scale fusion $\{\bm{X}_{n}\}$.
@@ 4.3 Ablation Study
We use SLATE on ClevrTex as an example to evaluate how different designs contribute to our MSF. We use ARI+ARIfg as the metrics because ARI is mostly dominated by background segmentation performance while ARIfg only measures the foreground. We employ naive CNN (Kipf et al., 2022) as the primary encoder to reduce experiment time.
@@ A.1 Extended Related Work
Multi-scale & image/feature pyramid. Existing multi-scale methods (i) Either employ an image pyramid (Adelson et al., 1984; Singh & Davis, 2018; Najibi et al., 2019) without information fusion among pyramid levels, wasting useful information among different scales; (ii) Or rely on feature pyramids, namely, FPN and numerous variants (Lin et al., 2017; Tan et al., 2020; Piao et al., 2021), where features of different DNN depths are fused by channel-concat or element-wise-sum, mixing both low- and high-quality representations together. In contrast, our MSF is the first to enable fusion of multiple scales on VAE representations. By leveraging codebook matching (Van Den Oord et al., 2017), we selectively fuse high-quality information among scales, rather than mixing them together. As shown in Tab. 6, our MSF is superior to naive channel-concat or element-wise-sum.
@@
Multi-scale & VAE. Recent works like VAR (Tian et al., 2024) and SPAE (Yu et al., 2024) also build the multi-scale representations upon VAE. But they are different from our MSF as follows. (i) VAR vs MSF: VAR auto-regresses from smaller scales of VAE representation to the larger, but with no information fusion among scales. In contrast, our MSF realizes fusion among different scales on VAE for the first time. (ii) SPAE vs MSF: SPAE relies on multi-modal foundation model CLIP, while our MSF does not. SPAE element-wisely averages multiple scales into one, mixing both low- and high-quality representations together, i.e., no fusion among scales. Our MSF augments any scaled representation with all other scales’ high-quality representations, i.e., fusion among scales.
Relationships among SSLseg, OCL and WMs. As shown in Fig. 4, (1) SSLseg (Self-Supervised Learning segmentation), e.g., HCL (Ge et al., 2023) and VideoCutLER (Wang et al., 2024), focuses on extracting segmentation masks; (2) OCL (Object-Centric Learning), e.g., SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b), represents each object as a feature vector, with segmentation masks as byproducts, kind of overlapping with SSLseg; (3) WMs (World Models), e.g., SlotFormer (Wu et al., 2023a), upon OCL, address downstream tasks like visual reasoning, planning and decision-making.
@@ A.2 Extended Experiments
MSF on OpenImages subset with higher resolution. We use input resolutions 128 and 256 to evaluate on OpenImages subset Ge et al. (2023). Under resolution 128, we use $n$=3 and 4; under resolution 256, we use $n$=3, 4 and 5. Results are shown below Tab. 8. We use ARI+ARIfg as the metrics because ARI mostly reflect how well the background is segmented and ARIfg only measures foreground objects. According to the results, under resolution 128, n=3 is the best choice; under resolution 256, n=4 is the best choice.
resolution 128$\times$128 256$\times$256
@@ A.3 Model Architecture
To save computation while maintain strong performance for all methods, we adopt the big-little design used in (Zhao et al., 2024), which takes small resolution 128 as input while still produce high resolution (128) feature maps. We employ the DINO v1, tiny, with a downsampling rate of 8, as presented in (Caron et al., 2021). We integrate this into a big-little architecture by running DINO in parallel with the previously mentioned naive CNN. The output of DINO is then projected, upsampled, and combined element-wise with the output of the CNN. The architecture is briefly represented as dino()-conv(k3s1p1)-upsample(k8) + naivecnn(). Notably, the weights of DINO are frozen, which retains its pretrained features while allowing the CNN to focus on refining high spatial details.
@@
We adhere to the standard design outlined in (Locatello et al., 2020). This design is consistent across all experiments utilizing this module.
@@
For transformer-based OCL, e.g., SLATE and STEVE, we follow the architecture described in (Singh et al., 2022a), unified across all experiments involving this component. This decoder operates on VAE intermediate feature tensors, which undergo 4x4 spatial downsampling relative to the input. SlotAttention-derived slots are provided as the conditioning input. The input channel dimension for this module is set to 256.
For diffusion-based OCL, e.g., SlotDiffusion, we follow the architecture described in (Wu et al., 2023b). This decoder take VAE representation with noise as input, taking slots, time step as condition, finally produce the reconstruction of the noise being added to the input. The input channel dimension for this module is set to 4.
@@
The architecture of this module is in accordance with the design described in (Singh et al., 2022a), and it is unified for all experiments incorporating this module. The VAE encoder applies a 4x4 spatial downsampling, and the VAE decoder performs a corresponding 4x4 upsampling.
@@
For the non-grouped codebook design, we adhere to the typical dVAE and VQ-VAE implementations as described in (Singh et al., 2022a) and (Wu et al., 2023b), respectively. For the grouped codebook design used in SLATE/STEVE and SlotDiffusion, we take the GDR as the basis, by applying with our design as mentioned in main body of this paper. Thus except what we describe in the main body, we also adopt annealing residual connection and Gumbel sampling code matching to stabilize training.

### v3

@@ 4.2 VAE Guidance
To investigate how MSF fuses multi-scale information into VAE guidance, we analyze object separability (stanic2024cae; lowe2024rvae). Empirically, better object separability in VAE guidance contributes to better OCL performance. We cluster on the VAE guidance $\bm{X}_{n=1}$ then calculate inter-/intra-cluster distances; We also visualize (caron2021dino) object separability of scale-invariant representations $\{\dot{\bm{X}}_{n}^{\mathrm{si}}\}$ and scale-variant representations $\{\bm{X}_{n}^{\mathrm{sv}}\}$, as well as representations before inter-scale fusion $\{\bm{X}_{n}^{\mathrm{si}}\}$ and representations after intra-scale fusion $\{\bm{X}_{n}\}$.
@@ 4.3 Ablation Study
We use SLATE on ClevrTex as an example to evaluate how different designs contribute to our MSF. We use ARI+ARIfg as the metrics because ARI is mostly dominated by background segmentation performance while ARIfg only measures the foreground. We employ naive CNN (kipf2021savi) as the primary encoder to reduce experiment time.
@@ A.1 Extended Related Work
Multi-scale & image/feature pyramid. Existing multi-scale methods (i) Either employ an image pyramid (adelson1984imagepyramid; singh2018snip; najibi2019sniper) without information fusion among pyramid levels, wasting useful information among different scales; (ii) Or rely on feature pyramids, namely, FPN and numerous variants (lin2017featurepyramid; tan2020efficientdet; piao2021panet), where features of different DNN depths are fused by channel-concat or element-wise-sum, mixing both low- and high-quality representations together. In contrast, our MSF is the first to enable fusion of multiple scales on VAE representations. By leveraging codebook matching (van2017vqvae), we selectively fuse high-quality information among scales, rather than mixing them together. As shown in Tab. 6, our MSF is superior to naive channel-concat or element-wise-sum.
@@
Multi-scale & VAE. Recent works like VAR (tian2024var) and SPAE (yu2024spae) also build the multi-scale representations upon VAE. But they are different from our MSF as follows. (i) VAR vs MSF: VAR auto-regresses from smaller scales of VAE representation to the larger, but with no information fusion among scales. In contrast, our MSF realizes fusion among different scales on VAE for the first time. (ii) SPAE vs MSF: SPAE relies on multi-modal foundation model CLIP, while our MSF does not. SPAE element-wisely averages multiple scales into one, mixing both low- and high-quality representations together, i.e., no fusion among scales. Our MSF augments any scaled representation with all other scales’ high-quality representations, i.e., fusion among scales.
Relationships among SSLseg, OCL and WMs. As shown in Fig. 4, (1) SSLseg (Self-Supervised Learning segmentation), e.g., HCL (ge2023hcl) and VideoCutLER (wang2024videocutler), focuses on extracting segmentation masks; (2) OCL (Object-Centric Learning), e.g., SLATE (singh2021slate) and SlotDiffusion (wu2023slotdiffuz), represents each object as a feature vector, with segmentation masks as byproducts, kind of overlapping with SSLseg; (3) WMs (World Models), e.g., SlotFormer (wu2022slotformer), upon OCL, address downstream tasks like visual reasoning, planning and decision-making.
@@ A.2 Extended Experiments
MSF on OpenImages subset with higher resolution. We use input resolutions 128 and 256 to evaluate on OpenImages subset ge2023hcl. Under resolution 128, we use nn=3 and 4; under resolution 256, we use nn=3, 4 and 5. Results are shown below Tab. 8. We use ARI+ARIfg as the metrics because ARI mostly reflect how well the background is segmented and ARIfg only measures foreground objects. According to the results, under resolution 128, n=3 is the best choice; under resolution 256, n=4 is the best choice.
resolution $\times128$ $\times256$
@@ A.3 Model Architecture
To save computation while maintain strong performance for all methods, we adopt the big-little design used in (zhao2024gdr), which takes small resolution 128 as input while still produce high resolution (128) feature maps. We employ the DINO v1, tiny, with a downsampling rate of 8, as presented in (caron2021dino). We integrate this into a big-little architecture by running DINO in parallel with the previously mentioned naive CNN. The output of DINO is then projected, upsampled, and combined element-wise with the output of the CNN. The architecture is briefly represented as dino()-conv(k3s1p1)-upsample(k8) + naivecnn(). Notably, the weights of DINO are frozen, which retains its pretrained features while allowing the CNN to focus on refining high spatial details.
@@
We adhere to the standard design outlined in (locatello2020slotattent). This design is consistent across all experiments utilizing this module.
@@
For transformer-based OCL, e.g., SLATE and STEVE, we follow the architecture described in (singh2021slate), unified across all experiments involving this component. This decoder operates on VAE intermediate feature tensors, which undergo 4x4 spatial downsampling relative to the input. SlotAttention-derived slots are provided as the conditioning input. The input channel dimension for this module is set to 256.
For diffusion-based OCL, e.g., SlotDiffusion, we follow the architecture described in (wu2023slotdiffuz). This decoder take VAE representation with noise as input, taking slots, time step as condition, finally produce the reconstruction of the noise being added to the input. The input channel dimension for this module is set to 4.
@@
The architecture of this module is in accordance with the design described in (singh2021slate), and it is unified for all experiments incorporating this module. The VAE encoder applies a 4x4 spatial downsampling, and the VAE decoder performs a corresponding 4x4 upsampling.
@@
For the non-grouped codebook design, we adhere to the typical dVAE and VQ-VAE implementations as described in (singh2021slate) and (wu2023slotdiffuz), respectively. For the grouped codebook design used in SLATE/STEVE and SlotDiffusion, we take the GDR as the basis, by applying with our design as mentioned in main body of this paper. Thus except what we describe in the main body, we also adopt annealing residual connection and Gumbel sampling code matching to stabilize training.